### v1

# Quantum Cosmology, Decoherence, and the Emergence of Classical Spacetime

Abstract
We analyze the emergence of classical cosmological spacetimes in quantum cosmology by computing the reduced density matrix for long-wavelength curvature perturbations. Starting from standard Hartle–Hawking and tunneling boundary conditions, we emphasize that semiclassical WKB structure and inflationary squeezing do not by themselves yield classicality. Tracing over unobserved degrees of freedom and using the influence functional formalism, we derive the decoherence functional for superhorizon curvature modes during inflation. For a light massive environmental scalar field in the Bunch–Davies vacuum, we obtain an explicit noise kernel and show how a nonzero mass regulates the infrared behavior. We then evaluate decoherence under horizon-based and EFT-motivated coarse grainings, finding efficient suppression of interference between macroscopically distinct perturbation histories in both cases. The analysis clarifies the distinct roles of boundary conditions (branch amplitudes) and decoherence (classical branch selection) and yields an emergent cosmological arrow of time through environment-induced entanglement.
Quantum cosmology, Decoherence, Influence functional, Inflation, Emergence of classicality
I Quantum Cosmology Before Classicality
I.1 Hilbert’s Sixth Problem and Emergent Classicality
Hilbert’s Sixth Problem calls for the axiomatization of physical theories [1] and, in particular, for a precise understanding of how macroscopic classical laws emerge from underlying microscopic dynamics. In modern language, this problem encompasses the emergence of irreversible, probabilistic, and effectively classical behavior from fundamentally quantum and reversible laws. While the problem was originally posed in the context of mechanics and probability theory, its scope naturally extends to quantum field theory, statistical mechanics, and cosmology.
In quantum cosmology, Hilbert’s Sixth Problem takes on a sharpened form: how do classical spacetime, classical matter fields, and classical stochastic perturbations arise from a quantum theory of gravity and quantum fields in curved spacetime? The Wheeler–DeWitt equation and its associated boundary conditions define a universal quantum state, but they do not by themselves explain why observations are accurately described by classical cosmology.
Recent formulations of Hilbert’s Sixth Problem emphasize the role of coarse graining, statistical limits, and emergent descriptions in bridging microscopic and macroscopic theories. In this spirit, the influence functional and decoherence framework provides a concrete and calculable realization of Hilbert’s program in cosmology. By explicitly tracing over unobserved degrees of freedom, one obtains effective dynamics for relevant variables that are irreversible, stochastic, and classical in precisely the sense required by Hilbert’s Sixth Problem.
The approach adopted in this work is directly motivated by these ideas and by recent analyses emphasizing Hilbert’s Sixth Problem as a guiding principle for foundational questions in physics [13]. While Ref. [13] concerns classical limits in kinetic theory, we cite it here as a representative modern implementation of Hilbert VI reasoning—deriving irreversible macroscopic equations from reversible microdynamics—and we apply the same logic to the quantum-to-classical transition in cosmology via coarse graining and decoherence. In particular, we build on the viewpoint that decoherence and coarse graining are not interpretational add-ons, but essential structural elements required to derive classical laws from quantum dynamics. The results presented below implement this program explicitly for cosmological perturbations during inflation, thereby providing a concrete cosmological realization of Hilbert’s Sixth Problem.
Quantum cosmology seeks to describe the universe by a wavefunctional
$\Psi[h_{ij}(\mathbf{x}),\Phi(\mathbf{x})],$ (1)
defined over three-geometries and matter field configurations. Two canonical proposals for specifying this wavefunction are the no-boundary proposal of Hartle and Hawking and the tunneling proposal of Vilenkin [2, 3]. In practice, both are implemented in a truncated configuration space consisting of minisuperspace variables, such as the scale factor $a$, together with quantum perturbations.
In the semiclassical regime, both proposals admit WKB-type expansions,
$\Psi[a,\zeta,\sigma]\simeq\sum_{\alpha}A_{\alpha}(a)\,\mathrm{e}^{\mathrm{i}S_{\alpha}(a)}\,\psi_{\alpha}[\zeta,\sigma;a],$ (2)
where $\alpha$ labels distinct semiclassical branches. While this structure allows approximate classical trajectories for $a(t)$ and approximate Schrödinger evolution for perturbations along each branch, the total state remains a coherent superposition of macroscopically distinct geometries. WKB behavior alone therefore does not explain the emergence of classical spacetime, since the universal wavefunctional in Eq. (1) remains a coherent superposition of macroscopically distinct geometries.
Inflationary dynamics leads to strong squeezing of cosmological perturbations, making correlators appear classical. However, squeezing does not suppress phase coherence or select a preferred basis. The quantum state remains pure, and interference between macroscopically distinct perturbations persists. Classicality therefore requires a further mechanism.
Observations necessarily access only a subset of degrees of freedom. The physically relevant object is thus the reduced density matrix obtained by tracing over unobserved variables, defined explicitly in Eq. (3). Classical spacetime can emerge only if this reduced density matrix becomes approximately diagonal in an appropriate basis [4, 5, 8].
II Reduced Density Matrix and Influence Functional
Figure 1: Schematic roadmap (adapted from standard decoherence frameworks, e.g. influence-functional constructions of the Feynman–Vernon type). Quantum-cosmological boundary conditions define the universal state $\Psi$. Tracing over unobserved degrees of freedom yields a reduced density matrix for long-wavelength perturbations. The influence functional encodes dissipation and noise; the associated decoherence functional suppresses interference between macroscopically distinct histories. Classical stochastic descriptions and an operational arrow of time emerge after coarse graining, with the detailed rate depending only weakly on the coarse-graining prescription.
We separate the degrees of freedom into long-wavelength curvature perturbations $\zeta_{L}$ (the system) and unobserved degrees of freedom, including short-wavelength modes $\zeta_{S}$ and additional fields $\sigma$ (the environment). The reduced density matrix is
$\rho_{L}[a,\zeta_{L};a^{\prime},\zeta_{L}^{\prime}]=\int\mathcal{D}\zeta_{S}\,\mathcal{D}\sigma\;\Psi[a,\zeta_{L},\zeta_{S},\sigma]\,\Psi^{*}[a^{\prime},\zeta_{L}^{\prime},\zeta_{S},\sigma].$ (3)
The time evolution of $\rho_{L}$ can be written using the Schwinger–Keldysh formalism [6, 7]. Integrating out the environment defines an influence functional, whose general structure is encoded in the influence action Eq. (4) $\mathcal{F}=\mathrm{e}^{\mathrm{i}S_{\mathrm{IF}}}$, where the influence action admits the general form
$S_{\mathrm{IF}}=\int\mathrm{d}^{4}x\,\mathrm{d}^{4}x^{\prime}\left[\zeta_{\Delta}(x)D(x,x^{\prime})\zeta_{c}(x^{\prime})+\frac{\mathrm{i}}{2}\,\zeta_{\Delta}(x)N(x,x^{\prime})\zeta_{\Delta}(x^{\prime})\right]+\cdots.$ (4)
Here $D$ is a real, causal dissipation kernel and $N$ is a real, positive semidefinite noise kernel. Decoherence is governed by the imaginary term involving $N$.
The suppression of off-diagonal elements of the reduced density matrix is controlled by the decoherence functional defined in Eq. (5)
$\Gamma[\zeta_{\Delta}]=\frac{1}{2}\int\mathrm{d}^{4}x\,\mathrm{d}^{4}x^{\prime}\;\zeta_{\Delta}(x)N(x,x^{\prime})\zeta_{\Delta}(x^{\prime}).$ (5)
When $\Gamma\gg 1$ for macroscopically distinct configurations (Eq. (5)), the reduced density matrix becomes approximately diagonal.
III Explicit Noise Kernel for a Light Environment Field
We consider an environment consisting of a light scalar field $\sigma$ with mass $m_{\sigma}\lesssim H$ in the Bunch–Davies (BD) vacuum [9, 8]. The leading effective interaction with the long-wavelength curvature perturbation is modeled as
$S_{\mathrm{int}}=g\int\mathrm{d}\eta\,\mathrm{d}^{3}x\;a^{4}(\eta)\,\zeta_{L}(\eta,\mathbf{x})\,\sigma^{2}(\eta,\mathbf{x}),$ (6)
with $a(\eta)=-1/(H\eta)$. Such a coupling captures the universal fact that long-wavelength curvature perturbations modulate local volumes and hence the dynamics of additional sectors.
For a Gaussian environment state, the noise kernel appearing in Eq. (5) entering the influence action is [6, 7]
$N(x,x^{\prime})=g^{2}\,\mathrm{Re}\left[G^{>}_{\sigma}(x,x^{\prime})^{2}\right],$ (7)
where $G^{>}_{\sigma}(x,x^{\prime})=\langle\sigma(x)\sigma(x^{\prime})\rangle$ is the Wightman function. Using BD mode functions one finds
$G^{>}_{\sigma}(x,x^{\prime})=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\,\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot(\mathbf{x}-\mathbf{x}^{\prime})}u_{k}(\eta)u^{*}_{k}(\eta^{\prime}),$ (8)
with
$u_{k}(\eta)=\frac{\sqrt{\pi}}{2}e^{\mathrm{i}(\nu+1/2)\pi/2}H(-\eta)^{3/2}H^{(1)}_{\nu}(-k\eta),\qquad\nu=\sqrt{\frac{9}{4}-\frac{m_{\sigma}^{2}}{H^{2}}},$ (9)
where $H^{(1)}_{\nu}$ is a Hankel function. For a light field it is convenient to define
$\Delta=\frac{3}{2}-\nu\simeq\frac{m_{\sigma}^{2}}{3H^{2}},$ (10)
which controls the superhorizon time dependence.
Substituting Eqs. (8)–(9) into Eq. (7) yields the explicit convolution
$N(\eta,\eta^{\prime};\mathbf{k})=g^{2}\,\mathrm{Re}\left[\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}u_{q}(\eta)u^{*}_{q}(\eta^{\prime})\,u_{|\mathbf{k}-\mathbf{q}|}(\eta)u^{*}_{|\mathbf{k}-\mathbf{q}|}(\eta^{\prime})\right].$ (11)
For a superhorizon long mode $k_{L}\ll aH$, the integral is dominated by $q\gg k_{L}$ and simplifies to
$N(\eta,\eta^{\prime};k_{L})\simeq g^{2}\,\mathrm{Re}\left[\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\left(u_{q}(\eta)u^{*}_{q}(\eta^{\prime})\right)^{2}\right].$ (12)
In the superhorizon regime $q\eta\ll 1$, the mode functions scale as $u_{q}(\eta)\propto(-\eta)^{\Delta}q^{-3/2+\Delta}$, implying
$N(\eta,\eta^{\prime};k_{L})\propto g^{2}H^{4}(-\eta)^{2\Delta}(-\eta^{\prime})^{2\Delta}\int\mathrm{d}q\,q^{-4+4\Delta}.$ (13)
Thus a nonzero mass $m_{\sigma}\neq 0$ softens the infrared behavior and renders the accumulation of decoherence physically controlled by the finite duration and size of the inflationary patch, rather than by a formal divergence. This explicit kernel provides the basis for the coarse-grained estimates discussed in the next section.
IV Coarse Graining and Emergent Classical Dynamics
In this section we evaluate the decoherence functional derived above under two physically motivated coarse-graining prescriptions, following the general logic developed in the cosmological decoherence literature [9, 11, 12]. While the precise numerical rate of decoherence depends on the coarse graining, its qualitative behavior—namely the efficient suppression of interference between macroscopically distinct curvature perturbations—is robust.
IV.1 Horizon-Based Coarse Graining
A natural prescription in inflationary cosmology is to define the environment as modes whose physical wavelength is shorter than the Hubble radius at a given time. Concretely, we split modes according to
$q>k_{c}(\eta)\equiv\epsilon\,a(\eta)H,$ (14)
where $\epsilon\ll 1$ is a fixed numerical parameter. Modes continuously cross from the environment into the system as inflation proceeds, but decoherence is generated primarily before horizon crossing, when environmental modes are still oscillatory.
For subhorizon modes $q\gg aH$, the BD mode functions admit a WKB approximation,
$u_{q}(\eta)\simeq\frac{1}{\sqrt{2q}}\exp(-\mathrm{i}q\eta),$ (15)
so that
$\left(u_{q}(\eta)u_{q}^{*}(\eta^{\prime})\right)^{2}\simeq\frac{1}{4q^{2}}\exp[-2\mathrm{i}q(\eta-\eta^{\prime})].$ (16)
Substituting into the noise kernel yields
$N(\eta,\eta^{\prime};k_{L})\simeq\frac{g^{2}}{4}\int_{q>k_{c}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\exp[-2\mathrm{i}q(\eta-\eta^{\prime})]}{q^{2}}.$ (17)
The rapid oscillations render the kernel sharply peaked near $\eta=\eta^{\prime}$, allowing the local approximation
$N(\eta,\eta^{\prime})\simeq\mathcal{N}(\eta)\,\delta(\eta-\eta^{\prime}).$ (18)
With this approximation, the decoherence functional for a frozen superhorizon mode becomes, upon using Eq. (5),
$\Gamma_{k_{L}}\simeq\frac{1}{2}|\Delta\zeta_{k_{L}}|^{2}\int\mathrm{d}\eta\,a^{4}(\eta)\mathcal{N}(\eta).$ (19)
Changing variables to the number of e-folds $N=\ln a$ gives the parametric estimate
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}N}\sim\frac{g^{2}}{H^{2}}\,|\Delta\zeta_{k_{L}}|^{2}\,a^{3}(N),$ (20)
up to numerical factors of order unity. Decoherence therefore accumulates monotonically during inflation, and once the condition $\Gamma_{k_{L}}\gg 1$ is satisfied (cf. Eq. (5)), $\Gamma_{k_{L}}\gg 1$ the reduced density matrix becomes effectively diagonal.
IV.2 EFT-Based Coarse Graining
An alternative prescription is motivated by effective field theory. One introduces a fixed physical cutoff $\Lambda_{\mathrm{phys}}$ satisfying $H\ll\Lambda_{\mathrm{phys}}\ll M_{\mathrm{Pl}}$, and defines the environment as modes with physical momenta above this scale,
$q>a(\eta)\Lambda_{\mathrm{phys}}.$ (21)
Modes traced out in this way never re-enter the system, rendering decoherence effectively irreversible.
For all such modes one has $q/a\gg H$, and the WKB approximation remains valid throughout inflation. The noise kernel again becomes local in time,
$N(\eta,\eta^{\prime})\simeq\mathcal{N}_{\Lambda}\delta(\eta-\eta^{\prime}),$ (22)
with
$\mathcal{N}_{\Lambda}\sim g^{2}\int_{q>a\Lambda_{\mathrm{phys}}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{1}{q^{2}}\sim g^{2}\Lambda_{\mathrm{phys}}a(\eta).$ (23)
The resulting decoherence functional scales as a direct consequence of Eq. (5),
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}N}\sim\frac{g^{2}\Lambda_{\mathrm{phys}}}{H}\ |\Delta\zeta_{k_{L}}|^{2}\ a^{4}(N).$ (24)
Decoherence is thus even more efficient than in the horizon-based prescription and remains insensitive to the details of ultraviolet completion.
In both coarse-graining schemes the essential conclusion is the same: interactions with unobserved degrees of freedom rapidly suppress interference between distinct long-wavelength curvature perturbations, providing a robust dynamical origin for classicality.
IV.3 Order-of-Magnitude Estimates
It is useful to translate the above scaling relations into simple numerical estimates. Observations of the cosmic microwave background imply a dimensionless scalar amplitude $A_{s}\simeq 2.1\times 10^{-9}$, corresponding to a typical root-mean-square curvature perturbation $\zeta_{\rm rms}\sim\sqrt{A_{s}}\sim 5\times 10^{-5}$. Macroscopically distinct semiclassical branches may therefore be characterized by differences $|\Delta\zeta_{k_{L}}|\sim n\,\zeta_{\rm rms}$ with $n\sim 10$100100.
Using the EFT-based coarse graining, the parametric estimate for the decoherence rate yields
$\Gamma_{k_{L}}(\Delta N)\sim\frac{g^{2}}{4}\,\frac{\Lambda_{\mathrm{phys}}}{H}\,|\Delta\zeta_{k_{L}}|^{2}\,\bigl(e^{4\Delta N}-1\bigr),$ (25)
where $\Delta N$ is the number of $e$-folds since horizon exit of the mode $k_{L}$. Taking a representative ratio $\Lambda_{\mathrm{phys}}/H\sim 10$, one finds that decoherence, defined by $\Gamma_{k_{L}}\gtrsim 1$, typically occurs within a few to at most $\mathcal{O}(10)$ $e$-folds for couplings as small as $g\sim 10^{-5}$ and much more rapidly for matter-sector or portal-like couplings $g\sim 10^{-3}$$10^{-1}$. These estimates indicate that environment-induced decoherence is efficient well before the end of inflation for a broad and physically reasonable range of parameters.
V Hartle–Hawking and Tunneling Proposals After Decoherence
We now return to the interpretation of quantum-cosmological boundary conditions in light of the decoherence mechanism developed above. Both the no-boundary proposal of Hartle and Hawking and the tunneling proposal of Vilenkin define pure quantum states of the universe and, by themselves, do not yield classical spacetimes. Decoherence provides the missing dynamical ingredient.
V.1 Semiclassical Branch Structure
In the semiclassical regime, the universal wavefunction admits a WKB decomposition of the form
$\Psi[a,\zeta,\sigma]\simeq\sum_{\alpha}A_{\alpha}(a)\exp\left(\mathrm{i}S_{\alpha}(a)\right)\psi_{\alpha}[\zeta,\sigma;a],$ (26)
where the label $\alpha$ distinguishes different semiclassical branches. For homogeneous minisuperspace variables these branches typically correspond to expanding and contracting classical solutions, while for perturbations they encode different inflationary histories. Importantly, the Wheeler–DeWitt equation is time-reversal invariant, and both types of branches generically appear in the wavefunction for either boundary condition [2, 3].
At the level of the full wavefunction, these branches coexist coherently. The WKB approximation alone does not suppress interference between macroscopically distinct geometries, nor does it select a preferred classical history.
V.2 Reduced Density Matrix in the Branch Basis
The physically relevant object is the reduced density matrix obtained after tracing over unobserved degrees of freedom. Projecting onto the semiclassical branch basis, one finds matrix elements of the schematic form
$\rho_{\alpha\beta}\equiv\int\mathcal{D}\zeta_{S}\mathcal{D}\sigma\Psi_{\alpha}\Psi^{*}_{\beta}\propto A_{\alpha}A^{*}_{\beta}\exp\left[-\Gamma_{\alpha\beta}\right],$ (27)
where $\Gamma_{\alpha\beta}$ is the decoherence functional evaluated for the difference between the long-wavelength field configurations associated with branches $\alpha$ and $\beta$.
When $\Gamma_{\alpha\beta}\gg 1$, the off-diagonal elements of the reduced density matrix are exponentially suppressed, and the density matrix becomes approximately diagonal in the branch index. Only at this stage does it become meaningful to interpret the wavefunction as describing an ensemble of mutually exclusive classical histories.
V.3 Expanding versus Contracting Histories
A particularly important application concerns the distinction between expanding and contracting semiclassical branches. In an expanding universe, the available phase space of environmental degrees of freedom grows rapidly with time, leading to strong entanglement between the system and the environment. As shown in Sec. IV, this results in efficient and effectively irreversible decoherence.
In contrast, contracting branches generally generate less entanglement with environmental degrees of freedom. As a result, interference between expanding and contracting branches is suppressed primarily in favor of the expanding histories, which acquire robust classical records. This dynamical asymmetry gives operational content to the intuition often associated with the tunneling proposal, but it arises only after decoherence is taken into account.
V.4 Boundary Conditions versus Classicality
The analysis above clarifies the logically distinct roles played by boundary conditions and decoherence. The Hartle–Hawking and tunneling proposals determine the relative amplitudes $A_{\alpha}$ of semiclassical branches in the universal wavefunction. Decoherence, by contrast, determines which of these branches become classical by suppressing quantum interference.
Statements such as “the tunneling proposal predicts an expanding universe” are therefore incomplete unless supplemented by a discussion of environment-induced decoherence. Without decoherence, expanding and contracting branches remain part of a coherent superposition. With decoherence, expanding branches are dynamically selected as classical due to their stronger coupling to environmental degrees of freedom.
V.5 Emergent Arrow of Time
Neither the no-boundary nor the tunneling proposal imposes a fundamental arrow of time at the level of the Wheeler–DeWitt equation. In the present framework, the cosmological arrow of time emerges dynamically from the growth of entanglement between long-wavelength perturbations and their environment. Decoherence increases monotonically with expansion, rendering the arrow of time an emergent, rather than fundamental, property of classical cosmology.
VI Discussion and Outlook
We have shown that classical cosmology emerges dynamically through environment-induced decoherence. The stochastic description of inflationary perturbations arises as the late-time effective theory of a decohered quantum system [10]. Our results clarify the conceptual foundations of quantum cosmology and provide a concrete bridge between universal wavefunctions and classical spacetime histories.
Future directions include extending the analysis to reheating, tensor perturbations, and quantum-information-theoretic characterizations of cosmological decoherence.
VII Conclusion
In this work we have addressed the longstanding question of how classical spacetime and classical cosmological perturbations emerge from quantum cosmology. Starting from standard quantum-cosmological boundary conditions, including the no-boundary and tunneling proposals, we emphasized that semiclassical WKB behavior and squeezing of perturbations are not by themselves sufficient to explain classicality.
By explicitly tracing over unobserved degrees of freedom and employing the influence functional formalism, we derived a decoherence functional for long-wavelength curvature perturbations during inflation. We showed that interactions with environmental degrees of freedom lead to an efficient and robust suppression of quantum interference between macroscopically distinct perturbation histories. This mechanism renders the reduced density matrix approximately diagonal and justifies the use of classical stochastic descriptions in cosmology.
We demonstrated that the emergence of classicality is largely insensitive to the details of coarse graining, occurring both for horizon-based and effective-field-theory–motivated splits between system and environment. Classical cosmological histories therefore arise as a generic consequence of inflationary dynamics in the presence of additional degrees of freedom, rather than as a fine-tuned outcome of specific assumptions.
Our analysis clarifies the distinct roles played by quantum-cosmological boundary conditions and decoherence. Boundary conditions determine the relative amplitudes of semiclassical branches in the universal wavefunction, while decoherence determines which of these branches become classical. In particular, expanding cosmological histories are dynamically selected as classical due to their enhanced entanglement with environmental degrees of freedom, providing an emergent arrow of time without imposing one at the fundamental level.
Taken together, these results place the foundations of classical cosmology on the same footing as other emergent phenomena in quantum physics. They provide a coherent and quantitatively controlled framework connecting quantum cosmology, decoherence, and the classical universe described by cosmological observations.